PLEASE SHOW WORK STEP BY STEP THAN YOU

17. An asset promises to pay the following:

  • $60 each year for the next ten years: and
  • $1,000 in ten years

Assume all the cash flows are discounted by 6%. Use the annuity formula to get the price of the first part. Use the standard discounting formula to get the price of the second part. Add them together. This is a bond! It is described as paying a coupon rate of 60/1,000 = 6% (a coupon of 60), with a face value of 1,000 and a maturity of ten years. And its “yield-to-maturity” is 6%.

18/Optional. The previous question showed that combining an annuity (coupon) with a single cash flow (face value) produces a bond. Doing so gives the bond price formula:

P = coupon/yield + (face value – coupon/yield) / (1 + yield)maturity.  

  1. Using this formula, calculate the price of a 20-year maturity bond, face value 100 and coupon 5 for yield-to-maturity 4%, 5% & 6%
  2. Now perform a) for a 30-year bond with same coupon and face value.
Answer #1

Bond price is same as face value when the coupon rate is same as YTM, so if Coupon is 5% and Face value is 100 and YTM is 5% then bond price is 100. If the YTM > Coupon rate, then bond price is less than face value and if YTM < Coupon rate, then bond price is greater than face value. This is shown as below:

a) 20 year maturity bond:

Bond price for 20 year matuirty bond
Year Coupon (Year 1-20) of 5 and Face value Year 20 of 100 (Col A) YTM 4% (Discount rates) Col B Price of bond (Col C = Col A divided by Col B) YTM 5% (Discount rates) Col D Price of bond (Col E = Col A divided by Col D) YTM 5% (Discount rates) Col F Price of bond (Col G = Col A divided by Col F)
1 5 1.04^1=      1.0400                       4.81 1.05^1=      1.0500                      4.76 1.06^1=      1.0600                     4.72
2 5 1.04^2=      1.0816                       4.62 1.05^2=      1.1025                      4.54 1.06^2=      1.1236                     4.45
3 5 1.04^3=      1.1249                       4.44 1.05^3=      1.1576                      4.32 1.06^3=      1.1910                     4.20
4 5 1.04^4=      1.1699                       4.27 1.05^4=      1.2155                      4.11 1.06^4=      1.2625                     3.96
5 5 1.04^5=      1.2167                       4.11 1.05^5=      1.2763                      3.92 1.06^5=      1.3382                     3.74
6 5 1.04^6=      1.2653                       3.95 1.05^6=      1.3401                      3.73 1.06^6=      1.4185                     3.52
7 5 1.04^7=      1.3159                       3.80 1.05^7=      1.4071                      3.55 1.06^7=      1.5036                     3.33
8 5 1.04^8=      1.3686                       3.65 1.05^8=      1.4775                      3.38 1.06^8=      1.5938                     3.14
9 5 1.04^9=      1.4233                       3.51 1.05^9=      1.5513                      3.22 1.06^9=      1.6895                     2.96
10 5 1.04^10=      1.4802                       3.38 1.05^10=      1.6289                      3.07 1.06^10=      1.7908                     2.79
11 5 1.04^11=      1.5395                       3.25 1.05^11=      1.7103                      2.92 1.06^11=      1.8983                     2.63
12 5 1.04^12=      1.6010                       3.12 1.05^12=      1.7959                      2.78 1.06^12=      2.0122                     2.48
13 5 1.04^13=      1.6651                       3.00 1.05^13=      1.8856                      2.65 1.06^13=      2.1329                     2.34
14 5 1.04^14=      1.7317                       2.89 1.05^14=      1.9799                      2.53 1.06^14=      2.2609                     2.21
15 5 1.04^15=      1.8009                       2.78 1.05^15=      2.0789                      2.41 1.06^15=      2.3966                     2.09
16 5 1.04^16=      1.8730                       2.67 1.05^16=      2.1829                      2.29 1.06^16=      2.5404                     1.97
17 5 1.04^17=      1.9479                       2.57 1.05^17=      2.2920                      2.18 1.06^17=      2.6928                     1.86
18 5 1.04^18=      2.0258                       2.47 1.05^18=      2.4066                      2.08 1.06^18=      2.8543                     1.75
19 5 1.04^19=      2.1068                       2.37 1.05^19=      2.5270                      1.98 1.06^19=      3.0256                     1.65
20 5 1.04^20=      2.1911                       2.28 1.05^20=      2.6533                      1.88 1.06^20=      3.2071                     1.56
20 100 1.04^20=      2.1911                     45.64 1.05^20=      2.6533                    37.69 1.06^20=      3.2071                  31.18
Price of bond = Total of above 20 year cash flow                  113.59                 100.00                  88.53
At YTM 4% At YTM 5% At YTM 6%
Price of bond as calculated: Sum of all (Coupons/discounted rate) = 67.95 as table above 62.31 57.35
Face value (100/2.1911) = 45.64 37.69 31.18
Total of above = 67.95+45.64 = 113.59 100 88.53
b) Bond price for 30 year maturity
Year Coupon (Year 1-20) of 5 and Face value Year 20 of 100 (Col A) YTM 4% (Discount rates) Col B Price of bond (Col C = Col A divided by Col B) YTM 5% (Discount rates) Col D Price of bond (Col E = Col A divided by Col D) YTM 5% (Discount rates) Col F Price of bond (Col G = Col A divided by Col F)
1 5 1.04^1=      1.0400                       4.81 1.05^1=      1.0500                      4.76 1.06^1=      1.0600                     4.72
2 5 1.04^2=      1.0816                       4.62 1.05^2=      1.1025                      4.54 1.06^2=      1.1236                     4.45
3 5 1.04^3=      1.1249                       4.44 1.05^3=      1.1576                      4.32 1.06^3=      1.1910                     4.20
4 5 1.04^4=      1.1699                       4.27 1.05^4=      1.2155                      4.11 1.06^4=      1.2625                     3.96
5 5 1.04^5=      1.2167                       4.11 1.05^5=      1.2763                      3.92 1.06^5=      1.3382                     3.74
6 5 1.04^6=      1.2653                       3.95 1.05^6=      1.3401                      3.73 1.06^6=      1.4185                     3.52
7 5 1.04^7=      1.3159                       3.80 1.05^7=      1.4071                      3.55 1.06^7=      1.5036                     3.33
8 5 1.04^8=      1.3686                       3.65 1.05^8=      1.4775                      3.38 1.06^8=      1.5938                     3.14
9 5 1.04^9=      1.4233                       3.51 1.05^9=      1.5513                      3.22 1.06^9=      1.6895                     2.96
10 5 1.04^10=      1.4802                       3.38 1.05^10=      1.6289                      3.07 1.06^10=      1.7908                     2.79
11 5 1.04^11=      1.5395                       3.25 1.05^11=      1.7103                      2.92 1.06^11=      1.8983                     2.63
12 5 1.04^12=      1.6010                       3.12 1.05^12=      1.7959                      2.78 1.06^12=      2.0122                     2.48
13 5 1.04^13=      1.6651                       3.00 1.05^13=      1.8856                      2.65 1.06^13=      2.1329                     2.34
14 5 1.04^14=      1.7317                       2.89 1.05^14=      1.9799                      2.53 1.06^14=      2.2609                     2.21
15 5 1.04^15=      1.8009                       2.78 1.05^15=      2.0789                      2.41 1.06^15=      2.3966                     2.09
16 5 1.04^16=      1.8730                       2.67 1.05^16=      2.1829                      2.29 1.06^16=      2.5404                     1.97
17 5 1.04^17=      1.9479                       2.57 1.05^17=      2.2920                      2.18 1.06^17=      2.6928                     1.86
18 5 1.04^18=      2.0258                       2.47 1.05^18=      2.4066                      2.08 1.06^18=      2.8543                     1.75
19 5 1.04^19=      2.1068                       2.37 1.05^19=      2.5270                      1.98 1.06^19=      3.0256                     1.65
20 5 1.04^20=      2.1911                       2.28 1.05^20=      2.6533                      1.88 1.06^20=      3.2071                     1.56
21 5 1.04^21=      2.2788                       2.19 1.05^21=      2.7860                      1.79 1.06^21=      3.3996                     1.47
22 5 1.04^22=      2.3699                       2.11 1.05^22=      2.9253                      1.71 1.06^22=      3.6035                     1.39
23 5 1.04^23=      2.4647                       2.03 1.05^23=      3.0715                      1.63 1.06^23=      3.8197                     1.31
24 5 1.04^24=      2.5633                       1.95 1.05^24=      3.2251                      1.55 1.06^24=      4.0489                     1.23
25 5 1.04^25=      2.6658                       1.88 1.05^25=      3.3864                      1.48 1.06^25=      4.2919                     1.16
26 5 1.04^26=      2.7725                       1.80 1.05^26=      3.5557                      1.41 1.06^26=      4.5494                     1.10
27 5 1.04^27=      2.8834                       1.73 1.05^27=      3.7335                      1.34 1.06^27=      4.8223                     1.04
28 5 1.04^28=      2.9987                       1.67 1.05^28=      3.9201                      1.28 1.06^28=      5.1117                     0.98
29 5 1.04^29=      3.1187                       1.60 1.05^29=      4.1161                      1.21 1.06^29=      5.4184                     0.92
30 5 1.04^30=      3.2434                       1.54 1.05^30=      4.3219                      1.16 1.06^30=      5.7435                     0.87
30 100 1.04^30=      3.2434                     30.83 1.05^30=      4.3219                    23.14 1.06^30=      5.7435                  17.41
Price of bond = Total of above 30 year cash flow                  117.29                 100.00                  86.24
At YTM 4% At YTM 5% At YTM 6%
Price of bond as calculated: Sum of all (Coupons/discounted rate) = 86.46 as table above 76.86 68.83
Face value (100/3.2434) = 30.83 23.14 17.41
Total of above = 86.46+30.83 = 117.29 100 86.24